Two motorcycles start at the same place but travel in opposite directions. One car averages 32 miles per hour and the other averages 28 miles per hour. In how many hours will the two motorcycles be 150 miles apart?
Answer by checkley79(3341)

32X+28X=150
60X=150
X=150/60
X=2.5 HOURS THEY WILL BE 150 MILES APART.
PROOF:
32*2.5+28*2.5=150
80+70=150
150=150